# Tradeoffs Between Alignment and Helpfulness in Language Models with Representation Engineering

### v3

@@ Appendix B Proof of theorem 1
The theorem utilizes assumptions 1 and 2. The behavior expectation is:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]=\frac{\sum_{a_{+}\in good}\mathbb{P}_{\theta,r_{e}}(a_{+}|q)-\sum_{a_{-}\in bad}\mathbb{P}_{\theta,r_{e}}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta,r_{e}}(a_{+}|q)+\sum_{a_{-}\in bad}\mathbb{P}_{\theta,r_{e}}(a_{-}|q)}=$ (13)
@@
From $\Delta$ margin SVM classification of $\{U^{T}a_{+}\}_{a_{+}\in good}$ and $\{U^{T}a_{-}\}_{a_{-}\in good}$ by $\frac{\delta r(q)}{|\delta r(q)|}$ (assumption 2):
@@
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>\frac{1-\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\Delta|\delta r|)}}{1+\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\Delta|\delta r|)}}=$ (22)
$=\frac{1-\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)}{1+\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)}$ (23)
$=tanh(\frac{\Delta|\delta r|-\ln(\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)})}{2})$ (24)
@@
$\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}=\frac{1-B_{0}}{1+B_{0}}$ (25)
Where $B_{0}=B[\mathbb{P}_{\theta}(\cdot|q)]$, and that:
@@
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta|\delta r(q)|}{2}+arctanh(B_{0}))$ (27)
@@
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta\lambda}{2}\cdot r_{e}+arctanh(B_{0}))$ (28)
@@ Appendix C Proof of theorem 2
The theorem utilizes assumptions 1 and 3. Notice that:
$\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)=\frac{\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)}{1}=\frac{\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)}{\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)+\sum_{i\in incorrect}\mathbb{P}_{\theta,r_{e}}(a_{i}|q)}=$ (29)
$=\frac{\mathbb{P}_{\theta}(a_{correct}|q)}{\mathbb{P}_{\theta}(a_{correct}|q)+\sum_{i\in incorrect}\mathbb{P}_{\theta}(a_{i}|q)e^{\langle\delta r_{e}(q),U^{T}(e_{i}-e_{correct}(q))\rangle}}\leq$ (30)
Denote $X_{i}=\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle$ and by $\mathbb{P}_{correct}^{0}$ the probability of answering correctly without representation engineering:
$=\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\sum_{i\in incorrect}\mathbb{P}_{\theta}(a_{i}|q)e^{|\delta r_{e}(q)|(X_{i}-X_{correct})}}\leq$ (31)
Denote by $I_{\pm}=\{i\in incorrect|\pm X_{i}-X_{correct}>0\}$. Using the AM-GM inequality:
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\sum_{i\in I_{+}}\mathbb{P}_{i}^{0}e^{\frac{\sum_{i\in I_{+}}\mathbb{P}_{i}^{0}(X_{i}-X_{correct})}{\sum_{i\in I_{+}}\mathbb{P}_{i}^{0}}|\delta r_{e}(q)|}+\sum_{i\in I_{-}}\mathbb{P}_{i}^{0}e^{\frac{\sum_{i\in I_{-}}\mathbb{P}_{i}^{0}(X_{i}-X_{correct})}{\sum_{i\in I_{-}}\mathbb{P}_{i}^{0}}|\delta r_{e}(q)|}}$ (32)
Denote by ±$\mathbb{P}_{\pm}=\sum_{i\in I\pm}\mathbb{P}_{i}^{0}$ and $c_{\pm}=\frac{\sum_{i\in I_{\pm}}\mathbb{P}_{i}^{0}(X_{i}-X_{correct})}{\sum_{i\in I_{\pm}}\mathbb{P}_{i}^{0}}$. We get:
$=\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\mathbb{P}_{+}e^{c_{+}|\delta r_{e}(q)|}+\mathbb{P}_{-}e^{c_{-}|\delta r_{e}(q)|}}$ (33)
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}(e^{c_{+}|\delta r_{e}(q)|}+e^{c_{-}|\delta r_{e}(q)|})}$ (34)
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}(1+\frac{1}{2}\min\{c_{-},c_{+}\}^{2}|\delta r_{e}(q)|^{2})}$ (35)
@@
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}(1+\frac{1}{2}\min\{c_{-},c_{+}\}^{2}\lambda^{2}|r_{e}|^{2})}$ (36)
Under the assumption that $X_{i}$ distribute normally (assumption 3), $c_{\pm}$ are a weighted sum of positive/negative half-normal distributions with parameter $\sigma$, which we can refactor to $\sigma^{2}\cdot c_{\pm}^{\prime}$ where $c_{\pm}^{\prime}$ are the same but with $\sigma^{\prime}=1$. Yielding:
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}(1+\frac{1}{2}\beta^{2}\sigma^{2}\lambda^{2}|r_{e}|^{2})}$ (37)
We denote $\alpha=\frac{\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}}{1-\mathbb{P}^{0}_{correct}}$.
Notice that with probability $\frac{1}{|V|}$ the set $I_{\pm}$ is empty, therefor with probability $1-\frac{2}{|V|}$ both sets are not empty, thus ±$\mathbb{P}_{\pm}>0$ and $c_{+}>0$, $c_{-}<0$.
Notice that $P(X_{correct}>X_{i})=\frac{1}{2}$, thus $i\in I_{+}$ with probability $\frac{1}{2}$. Therefor $\mathbb{P}_{+}$ is a weighted sum of Bernoulli variables with weights $\{\mathbb{P}^{0}_{i}\}_{i\in incorrect}$.
Appendix D Proof of Corollary 2 and 3
@@
$\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)<\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{|\delta r_{e}(q)|\frac{\sum_{i\in incorrect}\mathbb{P}^{0}_{i}(X_{i}-X_{correct})}{\sum_{i\in incorrect}\mathbb{P}^{0}_{i}}}}$ (38)
@@
$\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{c_{+}|\delta r_{e}(q)|}}=\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{c_{+}\lambda r_{e}}}$ (39)
@@
Similarly, with probability $\frac{1}{|V|}$, $I_{+}$ is empty, thus
$\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)<\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{c_{-}|\delta r_{e}(q)|}}=\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)<\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{c_{-}\lambda r_{e}}}$ (40)
@@
Thus for $r_{e}\rightarrow\infty$, with probability $1-\frac{2}{|V|}$, it is bounded by a term that approaches 00 (that of theorem 2), with probability $1/|V|$ another term that approaches 00 (the sigmoid with $c_{+}$), and with probability $1/|V|$ a term that approaches $1$ (the sigmoid with $c_{-}$). Hence the expectation value is bounded by $\frac{1}{|V|}$. This proves corollary 3.
For corollary 2, notice that with probability $1-\frac{2}{|V|}$, the helpfulness is bounded by the term in theorem 2, while with probability $\frac{1}{|V|}$ it is bounded by:
$\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{c_{+}|\delta r_{e}(q)|}}$ (41)
@@
$<\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})(1+c_{+}^{2}\lambda^{2}r_{e}^{2})}$ (42)
@@
$<\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})(1+c_{-}^{2}\lambda^{2}r_{e}^{2})}$ (43)
@@
$\mathbb{E}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]<(1-\frac{1}{|V|})\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\alpha(1-\mathbb{P}^{0}_{correct})(1+\min\{c_{-},c_{+},\beta\}^{2}\lambda^{2}r_{e}^{2})}+\frac{1}{|V|}$ (44)
Hence over an entire dataset, we take the minimum of $\min\{c_{-},c_{+},\beta\}$ for the different questions.
$\mathbb{E}_{q\in dataset}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]\leq(1-\frac{1}{|V|})\frac{1}{1+\alpha^{\prime}(1+\frac{\sigma^{2}\lambda^{2}\beta^{\prime 2}}{2}r_{e}^{2})}+\frac{1}{|V|}$ (45)
@@ E.1 Compute
All experiments were conducted using 2 NVIDIA A10 GPUs, 2 CPUs and 80GB of memory. Recreating all experiments takes in total no more than 1 hour.
@@ E.2 Obtaining the engineered representations
We use the code provided by Zou et al. (2023a) to perform representation engineering on the Llama 2 13B chat and Llama 2 13B models Touvron et al. (2023). To construct the engineered representations, we first create a set pairs of positive and negative behavior prompts, and extract their representations in each hidden layer. Then, the pairs of vectors are subtracted per layer. On these vector differences we perform PCA analysis to extract the highest order PCA vectors per layer as the representation engineered vectors, $\{r_{e}^{(l)}\}$.
We create behavior vectors of harmlessness and not-racist to align Llama 2 13b and harmfulness and racism vectors to misalign Llama 2 13b chat.
  • For creating racism vectors for Llama 2 13B chat, we use the StereoSet bias dataset Nadeem et al. (2020). Given a bias statement from the dataset, we apply it over the template of an aligned response: "[INST]QUESTION: Consider the statement: [STATEMENT] ANSWER: The statement is false and racist.[/INST]", and a template of a misaligned response: "[INST]QUESTION: Consider the statement: [STATEMENT] ANSWER: 100% [/INST]". Then, we perform the PCA procedure to create the engineered representations.
  • For the not-racist vectors for Llama 2 13b, we use the same dataset and template described above for the racist vectors of Llama 2 13b chat, only this time the desired behavior is opposite, so we subtract the representations of the prompts in reverse order.
  • For the harmfulness vectors for Llama 2 13B chat, we take harmful instructions from AdvBench dataset Robey et al. (2021; 2022) and unharmful instructions from shareGPT, and create pairs of prompts with the format "[INST][INSTRUCTION][/INST]", where the instruction is taken from one of the datasets (a pair consists of two prompts, one with a harmful instruction and one with an unharmful instruction).
  • For the harmless vectors for Llama 2 13B, we take harmful instructions from AdvBench dataset, create pairs of harmless and harmful responses to them, sampled from Llama2 13B chat. The created pairs of prompts follow the format: "[INST]Question:[INSTRUCTION] Answer: [RESPONSE(Llama2-13b-chat)][/INST]". The responses of Llama2 13b chat were sampled by injecting the model with harmfulness vectors to produce the misaligned answer, and the aligned answers were sampled without representation engineering.

### v4

@@ Appendix B Proof of theorem 1
The theorem utilizes assumptions 3 and 1. The behavior expectation is:
$B[P_{\theta,r_{e}}(\cdot|q)]=\frac{\sum_{a_{+}\in good}P_{\theta,r_{e}}(a_{+}|q)-\sum_{a_{-}\in bad}P_{\theta,r_{e}}(a_{-}|q)}{\sum_{a_{+}\in good}P_{\theta,r_{e}}(a_{+}|q)+\sum_{a_{-}\in bad}P_{\theta,r_{e}}(a_{-}|q)}=$ (12)
@@
From $\Delta$ margin linear classification of $\{U^{T}a_{+}\}_{a_{+}\in good}$ and $\{U^{T}a_{-}\}_{a_{-}\in good}$ by $\frac{\delta r(q)}{|\delta r(q)|}$ (assumption 1):
@@
$B[P_{\theta,r_{e}}(\cdot|q)]>\frac{1-\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\Delta|\delta r|)}}{1+\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\Delta|\delta r|)}}=$ (21)
$=\frac{1-\frac{\sum_{a_{-}\in bad}P_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}P_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)}{1+\frac{\sum_{a_{-}\in bad}P_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}P_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)}$ (22)
$=tanh(\frac{\Delta|\delta r|-\ln(\frac{\sum_{a_{-}\in bad}P_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}P_{\theta}(a_{+}|q)})}{2})$ (23)
@@
$\frac{\sum_{a_{-}\in bad}P_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}P_{\theta}(a_{+}|q)}=\frac{1-B_{0}}{1+B_{0}}$ (24)
Where $B_{0}=B[P_{\theta}(\cdot|q)]$, and that:
@@
$B[P_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta|\delta r(q)|}{2}+arctanh(B_{0}))$ (26)
@@
$B[P_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta\lambda}{2}\cdot r_{e}+arctanh(B_{0}))$ (27)
@@ Appendix C Proof of theorem 2
The theorem utilizes assumptions 3 and 2. Notice that:
$P_{\theta,r_{e}}(a_{correct}|q)=\frac{P_{\theta,r_{e}}(a_{correct}|q)}{1}=\frac{P_{\theta,r_{e}}(a_{correct}|q)}{P_{\theta,r_{e}}(a_{correct}|q)+\sum_{i\in incorrect}P_{\theta,r_{e}}(a_{i}|q)}=$ (28)
$=\frac{P_{\theta}(a_{correct}|q)}{P_{\theta}(a_{correct}|q)+\sum_{i\in incorrect}P_{\theta}(a_{i}|q)e^{\langle\delta r_{e}(q),U^{T}(e_{i}-e_{correct}(q))\rangle}}\leq$ (29)
Denote $X_{i}=\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle$ and by $P_{correct}^{0}$ the probability of answering correctly without representation engineering:
$=\frac{P^{0}_{correct}}{P^{0}_{correct}+\sum_{i\in incorrect}P_{\theta}(a_{i}|q)e^{|\delta r_{e}(q)|(X_{i}-X_{correct})}}\leq$ (30)
Next, by considering the sum only only over highest probability tokens making up $1-\epsilon$ of the probability mass, for which we denote the incorrect tokens sum as $incorrect(\epsilon)$:
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+\sum_{i\in incorrect(\epsilon)}P_{\theta}(a_{i}|q)e^{|\delta r_{e}(q)|(X_{i}-X_{correct})}}\leq$ (31)
Denote by $I_{\pm}=\{i\in incorrect(\epsilon)|\pm X_{i}-X_{correct}>0\}$. Using the AM-GM inequality:
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+\sum_{i\in I_{+}}P_{i}^{0}e^{\frac{\sum_{i\in I_{+}}P_{i}^{0}(X_{i}-X_{correct})}{\sum_{i\in I_{+}}P_{i}^{0}}|\delta r_{e}(q)|}+\sum_{i\in I_{-}}P_{i}^{0}e^{\frac{\sum_{i\in I_{-}}P_{i}^{0}(X_{i}-X_{correct})}{\sum_{i\in I_{-}}P_{i}^{0}}|\delta r_{e}(q)|}}$ (32)
Denote by $P_{\pm}=\sum_{i\in I\pm}P_{i}^{0}$ and $c_{\pm}=\frac{\sum_{i\in I_{\pm}}P_{i}^{0}(X_{i}-X_{correct})}{\sum_{i\in I_{\pm}}P_{i}^{0}}$. We get:
$=\frac{P^{0}_{correct}}{P^{0}_{correct}+P_{+}e^{c_{+}|\delta r_{e}(q)|}+P_{-}e^{c_{-}|\delta r_{e}(q)|}}$ (33)
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+\min\{P_{-},P_{+}\}(e^{c_{+}|\delta r_{e}(q)|}+e^{c_{-}|\delta r_{e}(q)|})}$ (34)
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+\min\{P_{-},P_{+}\}(1+\frac{1}{2}\min\{c_{-},c_{+}\}^{2}|\delta r_{e}(q)|^{2})}$ (35)
@@
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+\min\{P_{-},P_{+}\}(1+\frac{1}{2}\min\{c_{-},c_{+}\}^{2}\lambda^{2}|r_{e}|^{2})}$ (36)
Under the assumption that $X_{i}$ distribute randomly (assumption 2), $c_{\pm}$ are a weighted sum of positive/negative random variables with parameter $\sigma$, which we can refactor to $\sigma^{2}\cdot c_{\pm}^{\prime 2}$ where $c_{\pm}^{\prime}$ are the same but with $\sigma^{\prime}=1$. Yielding:
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+\min\{P_{-},P_{+}\}(1+\frac{1}{2}\beta^{2}\sigma^{2}\lambda^{2}|r_{e}|^{2})}$ (37)
We denote $\alpha=\frac{\min\{P_{-},P_{+}\}}{(1-P^{0}_{correct})(1-\epsilon)}$, since we considered only the tokens making $1-\epsilon$ of the probability mass, thus, $P_{+}+P_{-}=(1-\epsilon)(1-P_{correct}^{0})$. Hence $\alpha$ measures the non-tightness of the bound, due to the asymmetry between $P_{\pm}$, and $(1-\epsilon)$ the non-tightness due to not using all the words in the vocabulary for the bound, only the top $T$.
$=\frac{P^{0}_{correct}}{P^{0}_{correct}+(1-P_{correct}^{0})\alpha(1-\epsilon)(1+\frac{1}{2}\beta^{2}\sigma^{2}\lambda^{2}|r_{e}|^{2})}$ (38)
Notice that with probability $\frac{1}{T}$ the set $I_{\pm}$ is empty, therefor with probability $1-\frac{2}{T}$ both sets are not empty, thus $P_{\pm}>0$ and $c_{+}>0$, $c_{-}<0$.
Notice that $P(X_{correct}>X_{i})=\frac{1}{2}$, thus $i\in I_{+}$ with probability $\frac{1}{2}$. Therefor $P_{+}$ is a weighted sum of Bernoulli variables with weights $\{P^{0}_{i}\}_{i\in incorrect}$.
Appendix D Proof of Corollary 2
@@
$P_{\theta,r_{e}}(a_{correct}|q)<\frac{P^{0}_{correct}}{P^{0}_{correct}+(1-P^{0}_{correct})e^{|\delta r_{e}(q)|\frac{\sum_{i\in incorrect}P^{0}_{i}(X_{i}-X_{correct})}{\sum_{i\in incorrect}P^{0}_{i}}}}$ (39)
@@
$\frac{P^{0}_{correct}}{P^{0}_{correct}+(1-P^{0}_{correct})e^{c_{+}|\delta r_{e}(q)|}}=\frac{P^{0}_{correct}}{P^{0}_{correct}+(1-P^{0}_{correct})e^{c_{+}\lambda r_{e}}}$ (40)
@@
Similarly, with probability $\frac{1}{T}$, $I_{+}$ is empty, thus
$P_{\theta,r_{e}}(a_{correct}|q)<\frac{P^{0}_{correct}}{P^{0}_{correct}+(1-P^{0}_{correct})e^{c_{-}|\delta r_{e}(q)|}}=P_{\theta,r_{e}}(a_{correct}|q)<\frac{P^{0}_{correct}}{P^{0}_{correct}+(1-P^{0}_{correct})e^{c_{-}\lambda r_{e}}}$ (41)
@@
Thus for $r_{e}\rightarrow\infty$, with probability $1-\frac{2}{T}$, it is bounded by a term that approaches 00 (that of theorem 2), with probability $1/T$ another term that approaches 00 (the sigmoid with $c_{+}$), and with probability $1/T$ a term that approaches $1$ (the sigmoid with $c_{-}$). Hence the expectation value is bounded by $\frac{1}{T}$. This proves corollary 2.
For a combination of all these results, notice that with probability $1-\frac{2}{T}$, the helpfulness is bounded by the term in theorem 2, while with probability $\frac{1}{T}$ it is bounded by:
$\frac{P^{0}_{correct}}{P^{0}_{correct}+(1-P^{0}_{correct})e^{c_{+}|\delta r_{e}(q)|}}$ (42)
@@
$<\frac{P^{0}_{correct}}{P^{0}_{correct}+(1-P^{0}_{correct})(1+c_{+}^{2}\lambda^{2}r_{e}^{2})}$ (43)
@@
$<\frac{P^{0}_{correct}}{P^{0}_{correct}+(1-P^{0}_{correct})(1+c_{-}^{2}\lambda^{2}r_{e}^{2})}$ (44)
@@
$\mathbb{E}[P_{\theta,r_{e}}(a_{correct}|q)]<(1-\frac{1}{T})\frac{P^{0}_{correct}}{P^{0}_{correct}+\alpha(1-P^{0}_{correct})(1+\min\{c_{-},c_{+},\beta\}^{2}\lambda^{2}r_{e}^{2})}+\frac{1}{T}$ (45)
@@ E.1 Compute
All experiments were conducted using 2 NVIDIA A10 GPUs, 2 CPUs and 80GB of memory. Recreating all experiments takes in total no more than 4 hours.
@@ E.2 Obtaining the engineered representations
We use the code provided by Zou et al. (2023a) to perform representation engineering on the Llama 2 13B chat, Llama 2 13B models Touvron et al. (2023) and Llama 3.1 8B chat, Llama 3.1 8B models Dubey et al. (2024). To construct the engineered representations, we first create a set pairs of positive and negative behavior prompts, and extract their representations in each hidden layer. Then, the pairs of vectors are subtracted per layer. On these vector differences we perform PCA analysis to extract the highest order PCA vectors per layer as the representation engineered vectors, $\{r_{e}^{(l)}\}$.
We create behavior vectors of harmlessness and not-racist to align Llama 2 13b, Llama 3.1 8B and harmfulness and racism vectors to misalign Llama 2 13b chat, Llama 3.1 8B Instruct.
  • For creating racism vectors for Llama 2 13B chat and Llama 3.1 8B Instruct, we use the StereoSet bias dataset Nadeem et al. (2020). Given a bias statement from the dataset, we apply it over the template of an aligned response: “Qusestion: Consider the statement: [STATEMENT] Answer: The statement is false and racist.", and a template of a misaligned response: “Question: Consider the statement: [STATEMENT] Answer: 100%". Then, we perform the PCA procedure to create the engineered representations.
  • For the non-racist vectors for Llama 2 13B and Llama 3.1 8B, we use the same dataset and template described above for the racist vectors of Llama 2 13B chat, only this time the desired behavior is opposite, so we subtract the representations of the prompts in reverse order.
  • For the harmfulness vectors for Llama 2 13B chat and Llama 3.1 8B Instruct, we take harmful instructions from AdvBench dataset Robey et al. (2021; 2022) and unharmful instructions from shareGPT, and create pairs of prompts with the format “Request: Can you [INSTRUCTION]?", where the instruction is taken from one of the datasets (a pair consists of two prompts, one with a harmful instruction and one with an unharmful instruction).
  • For the harmless vectors for Llama 2 13B, we take harmful instructions from AdvBench dataset, create pairs of harmless and harmful responses to them, sampled from Llama2 13B chat. The created pairs of prompts follow the format: “[INST]Question:[INSTRUCTION] Answer: [RESPONSE(Llama2-13b-chat)][/INST]". The responses of Llama2 13b chat were sampled by injecting the model with harmfulness vectors to produce the misaligned answer, and the aligned answers were sampled without representation engineering. For Llama 3.1 8B, the usual method used for harmful vectors sufficed, only this time the desired behavior is opposite, so we subtract the representations of the prompts in reverse order.